# StoryTeller: Improving Long Video Description through Global Audio-Visual Character Identification

### v1

@@ Abstract
Existing large vision-language models (LVLMs) are largely limited to processing short, seconds-long videos and struggle with generating coherent descriptions for extended video spanning minutes or more. Long video description introduces new challenges, such as plot-level consistency across descriptions. To address these, we figure out audio-visual character identification, matching character names to each dialogue, as a key factor. We propose StoryTeller, a system for generating dense descriptions of long videos, incorporating both low-level visual concepts and high-level plot information. StoryTeller uses a multimodal large language model that integrates visual, audio, and text modalities to perform audio-visual character identification on minute-long video clips. The results are then fed into a LVLM to enhance consistency of video description. We validate our approach on movie description tasks and introduce MovieStory101, a dataset with dense descriptions for three-minute movie clips. To evaluate long video descriptions, we create MovieQA, a large set of multiple-choice questions for the MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments show that StoryTeller outperforms all open and closed-source baselines on MovieQA, achieving 9.5% higher accuracy than the strongest baseline, Gemini-1.5-pro, and demonstrating a +15.56% advantage in human side-by-side evaluations. Additionally, incorporating audio-visual character identification from StoryTeller improves the performance of all video description models, with Gemini-1.5-pro and GPT-4o showing relative improvement of 5.5% and 13.0%, respectively, in accuracy on MovieQA. Datasets and code are available at https://github.com/hyc2026/StoryTeller.
@@ 1 Introduction
Generating detailed video descriptions is a fundamental challenge in video understanding. While large vision-language models (LVLMs) have made significant progress [41, 24, 8, 40, 23, 22], they remain limited to processing seconds-long videos and generating low-level visual concepts like objects, scenes and atomic actions. Real-world videos, such as movies, are much longer and require high-level information for effective description, such as plot development. Recent research works address long videos by segmenting them, generating descriptions for each segment, and combining them using large language models [47, 13, 14]. However, there is still a gap to resolve the key challenge of long video description: maintaining consistency. This includes ensuring logical coherence in the storyline, consistency in character descriptions and motivations, and overall continuity and fluidity of the description.
Accurate character identification is essential for ensuring consistency in long video descriptions. While existing studies [32, 18] have improved the quality of video descriptions by focusing on character identification and tracking through visual elements, audio information is equally critical, as dialogue significantly contributes to drive the story progressing. Nevertheless, a gap remains in linking dialogue with on-screen characters to identify speakers, and further improve the capability for long video descriptions.
@@
A major challenge in long video description is the lack of training and test data. To address this, we introduce MovieStory101, a new dataset focused on movies—long videos with high information density. The videos in MovieStory101 are from Movie101[45] and Movie101v2 [46], comprising 187 movies split into 5,982 three-minute clips. Each clip is annotated with storyline descriptions, subtitles, and character identity labels for each dialogue line. The storyline annotations ensure a complete understanding of the movie’s plot. By training on this dataset, we aim to unleash the model’s ability to generate plot-level descriptions. In adiition, evaluating long video descriptions is also challenging. Besides human evaluation, we propose an automated evaluation method for MovieStory101’s test set: MovieQA. MovieQA has 38 multiple-choice questions in average for each video, covering visual actions, character relationships, and plot details. We evaluate a long video description by inputting it into GPT-4 [28] to answer these questions, using the accuracy as an evaluation metric.
Based on MovieQA and human side-by-side evaluation, we demonstrate that StoryTeller outperforms Gemini-1.5-pro [35], GPT-4o [29], and several advanced open-source LVLMs [24, 20, 41, 7] in generating descriptions for long videos. Specifically, the 7B model-based StoryTeller achieves an accuracy in the MovieQA evaluation that is 9.5% higher than the best-performing baseline model, Gemini-1.5-pro. In human side-by-side evaluation, StoryTeller shows an advantage of +15.56% over Gemini-1.5-pro. Additionally, we verify that incorporating audio-visual character identification results from StoryTeller as input features consistently enhances the long video description capabilities of various LVLMs. In particular, Gemini-1.5-pro and GPT-4o achieve relative improvements of 5.5% and 13.0% in accuracy on MovieQA, respectively.
@@ 2 Related Work
Long Video Description. The rapid advancement of LVLMs has greatly accelerated progress in video description tasks [41, 24, 8, 40, 23, 22]. However, most LVLMs are constrained by the limited number of frames they can process, making them effective primarily for short video clips spanning only a few seconds. These models typically generate descriptions focused on low-level visual concepts, such as objects, scenes, and actions. There remains a gap in describing longer videos, which may span several minutes or more, and in capturing higher-level content like character interactions and plot progression. To address this challenge, some research works [47, 13, 14] treat a long video as a series of short clips, incorporating the descriptions of previous clips when describing the current one. However, the quality of earlier segment descriptions significantly affects the accuracy of the current description, leading to cumulative errors that compromise the consistency and coherence of the overall narrative [13]. Another approach, Video ReCap [17], introduces a recursive video captioning model capable of handling videos of varying lengths. Nevertheless, for longer videos, this method tends to produce summaries, limiting its ability to generate detailed descriptions.
Audio-Visual Character Identification. Character identification is essential for generating high-quality video descriptions [32, 18], particularly to maintain consistency across different clips. Previous research has mainly focused on character identification at the visual level, identifying characters in individual frames [27, 38]. However, videos contain both audio and visual elements, and utilizing audio cues to identify which character is speaking can straightforwardly enhance video descriptions. A closely related area of research is audio-visual diarization, which integrates both visual and auditory features to segment speakers in audio [42, 15]. Nonetheless, current methods are still unable to identify the speakers’ identities.
@@ 3.1 Overview
StoryTeller is a fully automated system for generating descriptions of long videos, capable of processing videos that span several minutes. As illustrated in Figure 1, the system consists of three main modules. First, the video segmentation module divides the long video into multiple short clips, each lasting several seconds, while preserving the integrity and relative independence of these clips, as detailed in Appendix C. The second component is the audio-visual character identification module, which employs a dual mechanism, local and global, to accurately identify characters throughout the long video. Locally, a multimodal large language model (MLLM) integrates audio, visual, and text modalities to achieve character identification using both audio and visual cues within each short clip. Globally, during inference, a global decoding algorithm processes multiple short clips associated with the same character simultaneously, combining their local results from the MLLM to enhance the overall identification accuracy. Finally, the description generation module utilizes an LVLM to produce a detailed description. By incorporating the audio-visual character identification as input, the LVLM generates coherent video descriptions that span the entire long video, resulting in a plot-level dense narrative.
@@ 4 MovieStory101 Dataset
We introduce MovieStory101, a long video description dataset based on movies. We choose movies because they contain rich and diverse information. A high-quality movie description should capture not only fine-grained details such as scenes, characters, and actions but also higher-level aspects like character relationships, plot development, and causal links between events (see Figure. 3). Furthermore, since movies are inherently long videos, their descriptions require linguistic coherence and logical consistency throughout the narrative. Consequently, the task of movie description presents significant challenges.
4.1 Movie Source
We build MovieStory101 dataset based on the publicly available datasets, Movie101 [45] and Movie101v2 [46]. Together, these datasets include 203 movies, providing audio descriptions for scenes without dialogue in the barrier-free versions.
Since the movies in Movie101 and Movie101v2 contain audio descriptions integrated into the audio tracks, we collect the original editions of the movies to obtain the original audio tracks. After reviewing the content, we select 187 movies for our dataset. We exclude 16 movies because their original editions differed significantly from their barrier-free versions. Each selected movie is segmented into 3-minute clips. The final MovieStory101 dataset includes 5,350 / 140 / 492 clips in the training / development / testing split, totaling 300 hours. Each split contains video clips from different movies.
Type Question Optinos
@@ D. To charge her phone
Table 1: Examples of Action, Character, and Plot Questions in MovieQA.
@@ 4.2 Annotation Process
The annotation process comprises three steps: movie descriptions, character photos, and audio-visual character identification.
Movie description. We annotate each movie clip in the dataset with detailed description. To improve annotation efficiency, we present audio descriptions and ASR data from Movie101 and Movie101v2 as references to the annotators. Annotators are tasked with providing the descriptions of each event in the video, along with the corresponding start and end timestamps. The descriptions should include details about scenes, characters, actions, dialogues, and plot development. The aim is to create a coherent long-form description of the video. Figure 3 provides an example in MovieStory101 dataset. See annotation guidelines in Appendix A. In addition, annotators revise the ASR data in parallel to generate subtitles for each movie clip.
Character photos. For each movie clip, annotators identify main characters present and select a face for each character. A cast list is then created, to record each character’s name and corresponding face, as shown in Figure 3.
Audio-visual character identification. For each line of dialogue in the video clips, annotators identify the speaker. If the speaker is included in the cast list, the corresponding character name is assigned. Otherwise, a descriptive label is used to identify the speaker. Figure 3 provides an example of audio-visual character identification in the MovieStory101 dataset.
4.3 Evaluation of Movie Description
Evaluating long video descriptions is challenging due to the numerous details they contain, which increases evaluation costs. To address this, we develop an automatic evaluation method called MovieQA for the MovieStory101 test set. On average, each 3-minute video in MovieQA is accompanied by 38 multiple-choice questions focusing on visual actions, character relationships, and plot development. During evaluation, we employ GPT-4 [28] to answer these questions given the long description as context, and the QA accuracy serves as an automated metric for assessing the quality of the description. The prompt template is shown in Appendix F. Table 1 shows examples in MovieQA.
Creation for MovieQA We identify three key types of information that need to be highlighted in evaluation of long video descriptions: 1) Action: actions within the movie, including action recognition, the sequence of actions, and the causal relationships between them; 2) Character: character roles, their relationships with others, and how these relationships evolve; 3) Plot: causes and effects, character motivations, and event connections. Based on these categories, we generate question-answer pairs for each type. The detailed process is shown in Appendix B.
Model MovieQA Accuracy
@@ 4.3 Evaluation of Movie Description
Table 2: The results of video descriptions generated by StoryTeller and baselines evluated on MovieQA.
@@
5 Experiments
5.1 Baselines
We compare our method against two types of baselines:
For all baselines, we provide the video, cast list and subtitles as input. Specifically, for Gemini-1.5-pro, we directly input the entire 3-minute video, while for the other models, we first divide the video into 10-second clips. We then sample 8 frames from each clip as input to generate descriptions for the individual segments. Finally, we concatenate the descriptions from each clip to form a complete description of the entire 3-minute video.
5.2 Experimental Results
First, we evaluate the descriptions generated by our method and the baselines on the MovieQA dataset. As shown in Table 2, StoryTeller outperforms all baselines on MovieQA. Notably, StoryTeller achieves a 9.5% higher accuracy than the strongest baseline, Gemini-1.5-pro. Table 3 shows a video example featuring multiple characters, where StoryTeller effectively tracks their relationships, resulting in description with fewer hallucination compared to Gemini-15-pro. When analyzing performance across different question types in MovieQA, StoryTeller outperforms all baselines on action, character and plot questions, indicating that StoryTeller provides more accurate descriptions of both low-level visual details and high-level plot information.
@@ 5.2 Experimental Results
Table 5: Evaluation of StoryTeller as a general framework for improving video descriptions across vairous LVLMs. The performance on MovieQA when key components are incrementally added: video segmentation module, integration of the audio-visual character identification (with ‘Others’ label for unlisted characters), and the use of descriptive names for unlisted characters. The baseline approach involves segmenting videos into 10-second clips and generating descriptions without audio-visual character identification.
@@
The system design of StoryTeller is a general framework applicable to various LVLMs to enhance their long video descriptions. Table 5 presents the additive contributions of each component for various LVLMs. First, replacing the direct division into 10-second clips with our video segmentation module consistently improves description accuracy across all models, resulting in an average relative increase of 6.3% in accuracy on the MovieQA. Second, utilizing the audio-visual identification results generated by our model as input for the LVLMs, we observe an average relative accuracy improvement of 6.4%. This demonstrates that our model’s audio-visual identification results is effective for enhancing video descriptions across various LVLMs. Lastly, we observe that replacing the label ‘Others’ with descriptive names generated by our model further improves video descriptions, consistently benefiting all models. Notably, under our framework, Gemini-1.5-pro and GPT-4o achieve relative improvements of 5.5% and 13.0% in accuracy on MovieQA, respectively.
@@
At last, we conduct an ablation study to evaluate the necessity of fine-tuning for audio and visual tasks (audio diarization, character recognition, identification, and tracking), as well as effect of global decoding within the audio-visual identification module. We test the accuracy of audio-visual character identification on the MovieStory101 test set, with the results shown in Table 6. Notably, fine-tuning the audio and visual tasks improves performance by 4.6%, while global decoding contributes 3.2% improvement.

### v2

@@ Abstract
Existing large vision-language models (LVLMs) are largely limited to processing short, seconds-long videos and struggle with generating coherent descriptions for extended video spanning minutes or more. Long video description introduces new challenges, such as consistent character identification and plot-level descriptions incorporating both visual and audio information. To address these, we figure out audio-visual character identification, matching character names to each dialogue, as a key factor. We propose StoryTeller, a system for generating dense descriptions of long videos, incorporating both low-level visual concepts and high-level plot information. StoryTeller uses a multimodal large language model that integrates visual, audio, and text modalities to perform audio-visual character identification on minute-long video clips. The results are then fed into a LVLM to enhance consistency of video description. We validate our approach on movie description tasks and introduce MovieStory101, a dataset with dense descriptions for three-minute movie clips. To evaluate long video descriptions, we create StoryQA, a large set of multiple-choice questions for MovieStory101 test set. We assess descriptions by inputting them into GPT-4 to answer these questions, using accuracy as an automatic evaluation metric. Experiments show that StoryTeller outperforms all open and closed-source baselines on StoryQA, achieving 9.5% higher accuracy than the strongest baseline, Gemini-1.5-pro, and demonstrating a +15.56% advantage in human side-by-side evaluations. Additionally, incorporating audio-visual character identification from StoryTeller improves the performance of all video description models, with Gemini-1.5-pro and GPT-4o showing relative improvement of 5.5% and 13.0%, respectively, in accuracy on StoryQA. Datasets and code are available at https://github.com/hyc2026/StoryTeller.
@@ 1 Introduction
Generating detailed video descriptions is a fundamental challenge in video understanding. While large vision-language models (LVLMs) have made significant progress [45, 27, 8, 44, 26, 25], they remain limited to processing seconds-long videos and generating low-level visual concepts like objects, scenes and atomic actions. Real-world videos, such as movies, are much longer and require high-level information for effective description, such as plot development. Recent research works address long videos by segmenting them, generating descriptions for each segment, and combining them using large language models [51, 13, 14]. However, there is still a gap to resolve the key challenges of long video description: maintaining consistent character identification and incorporating dialogue, which conveys rich plot information.
Thus, accurate character identification and tracking across both visual and audio modalities are essential for high-quality long video descriptions. While existing studies [35, 20] have improved the quality of video descriptions by focusing on character identification and tracking through visual cues, a critical gap remains in linking dialogue with on-screen characters to identify speakers. Bridging this gap is essential for enhancing long video descriptions.
@@
A major challenge in long video description is the lack of training and test data. To address this, we introduce MovieStory101, a new dataset focused on movies—long videos with high information density. The videos in MovieStory101 are from Movie101[49] and Movie101v2 [50], comprising 187 movies split into 5,982 three-minute clips. Each clip is annotated with storyline descriptions, subtitles, and character identity labels for each dialogue line. The storyline annotations ensure a complete understanding of the movie’s plot. By training on this dataset, we aim to unleash the model’s ability to generate plot-level descriptions. In addition, evaluating long video descriptions is also challenging. Besides human evaluation, we propose an automated evaluation method for MovieStory101’s test set: StoryQA. StoryQA has 11 multiple-choice questions in average for each video, covering visual actions, character relationships, and plot details. We evaluate a long video description by inputting it into GPT-4 [31] to answer these questions, using the accuracy as an evaluation metric.
Based on StoryQA and human side-by-side evaluation, we demonstrate that StoryTeller outperforms Gemini-1.5-pro [38], GPT-4o [32], and several advanced open-source LVLMs [27, 23, 45, 7] in generating descriptions for long videos. Specifically, the 7B model-based StoryTeller achieves an accuracy in the StoryQA evaluation that is 9.5% higher than the best-performing baseline model, Gemini-1.5-pro. In human side-by-side evaluation, StoryTeller shows an advantage of +15.56% over Gemini-1.5-pro. Additionally, we verify the significant role of accurate audio-visual character identification on improving description. Incorporating audio-visual character identification results from StoryTeller as input consistently improves the performance of various LVLMs, with Gemini-1.5-pro and GPT-4o achieving relative accuracy gains of 5.5% and 13.0% on StoryQA, respectively.
@@ 2 Related Work
Long Video Description. Limited by the number of input frames, LVLMs are primarily effective for short video clips and typically generate descriptions focused on low-level visual concepts [45, 27, 8, 44, 26, 25]. There remains a gap in describing longer videos, which may span several minutes or more, and in capturing higher-level content like character interactions and plot progression.Some research [51, 13, 14] treats long videos as a series of short clips, incorporating previous descriptions when describing the current one, but this can lead to cumulative errors that affect consistency and coherence of the overall narrative [13]. Another approach, Video ReCap [19], offers a recursive model for varying video lengths. Nevertheless, for longer videos, this method tends to produce summaries, limiting its ability to generate detailed descriptions.
Audio-Visual Character Identification. Character identification is essential for high-quality video descriptions [35, 20], particularly to maintain consistency across clips. While prior research has focused on visual-level identification [30, 41], leveraging audio cues to determine which character is speaking can straightforwardly improve descriptions. A closely related area of research is audio-visual diarization, which integrates both visual and auditory features to segment speakers in audio [46, 16]. Nonetheless, current methods are still unable to identify the speakers’ identities.
@@ 3.1 Overview
StoryTeller is a fully automated system for generating descriptions of long videos. As shown in Figure 1, it consists of three main modules: 1) the video segmentation module divides the long video into multiple short clips, each lasting several seconds, while preserving the integrity and relative independence of these clips, as detailed in Appendix C; 2) the audio-visual character identification module, which employs a dual mechanism, local and global, to accurately identify characters throughout the long video. Locally, a multimodal large language model (MLLM) integrates audio, visual, and text modalities to achieve character identification using both audio and visual cues within each short clip. Globally, during inference, a global decoding algorithm processes multiple short clips associated with the same character simultaneously, combining their local results from the MLLM to enhance the overall identification accuracy; 3) the description generation module utilizes an LVLM to produce a detailed description. By incorporating the audio-visual character identification as input, the LVLM generates coherent video descriptions that span the entire long video, resulting in a plot-level dense narrative.
@@ 4 MovieStory101 Dataset
Dataset Video Source Description Style # Movie # Clip $L_{v}$ $L_{t}$ $N_{\text{char}}$
MovieNet [18] movie - 1100 42K 428.0 - -
ActivityNet Captions [22] ActivityNet activity description - 20k 180.0 49.2 -
YouCook2 [52] Youtube cook video activity description - 2k 316.2 67.8 -
ViTT [17] Youtube video activity tag - 8k 282.0 21.1 -
CMD-AD [15] movie audio description 1432 101k 2.8 9.1 -
Movie101 [49] movie audio description 101 14.1k 20.4 80.7 2.0
Movie101-v2 [50] movie audio description 203 46k 12.8 39.1 1.9
MovieStory101 movie full description 187 6k 180.0 565.0 4.3
Table 1: Data statistics for MovieStory101 compared with other video description datasets. $L_{v/t}$: average clip duration (seconds) or text length (words; Chinese characters underlined); $N_{\text{char}}$: average number of characters per clip.
Dataset
full
description
character
photo
subtitle
speaker
name
audio
description
MovieNet [18]
MovieQA [42]
CMD-AD [15]
Movie101 [49]
Movie101-v2 [50]
MovieStory101
Table 2: Annotation comparison between MovieStory101 and related movie datasets. Note that the subtitles provided by Movie101 and Movie101v2 are generated using ASR and MovieStory101 manually revised them.
We introduce MovieStory101, a long video description dataset based on movies. We choose movies because they contain rich and diverse information. A high-quality movie description should capture not only fine-grained details such as scenes, characters, and actions but also higher-level aspects like character relationships, plot development, and causal links between events (see Figure. 3).
The MovieStory101 task is partially related to audio description for movies [13, 14, 15], which narrates non-dialogue segments, and dense video captioning [22, 6], which generates temporally localized descriptions. However, MovieStory101 offers more comprehensive, plot-level description for long videos, integrating both visual and audio information. We refer to this as full description. In Table 1, we report the statistics of MovieStory101 and compare them with those of the other video description datasets, emphasizing its unique features: detailed descriptions, long video durations, and a large number of characters. Additionally, we provide a comparison of the annotations across different movie datasets in Table 2.
Type Question Options
@@ D. To charge her phone
Table 3: Examples of Action, Character, and Plot Questions in StoryQA.
4.1 Movie Source
We build MovieStory101 based on Movie101 [49] and Movie101v2 [50]. Together, they include 203 movies, providing audio descriptions for non-dialogue scenes.
Since the movies in Movie101 and Movie101v2 contain audio descriptions integrated into the audio tracks, we collect the original editions to obtain the original audio tracks. After review, we select 187 movies for our dataset, excluding 16 movies because their original editions differed significantly from their barrier-free versions. Each selected movie is segmented into 3-minute clips. The final MovieStory101 dataset includes 5,210 / 140 / 632 clips in the training / development / testing split, totaling 300 hours. Each split contains video clips from different movies.
Figure 4: Distribution of description length (left) and character count (right) in MovieStory101.
@@ 4.2 Annotation Process
The annotation process comprises three steps: movie descriptions, character photos, and audio-visual character identification. Figure 4 shows the statistics for descriptions and characters in MovieStory101.
Movie description. We annotate each movie clip with detailed description. To improve efficiency, annotators use audio descriptions and ASR data from Movie101 and Movie101v2 as references. Annotators are tasked with providing the descriptions of each event in the video, along with the corresponding start and end timestamps. The descriptions should include details about scenes, characters, actions, dialogues, and plot development. Figure 3 provides an example in MovieStory101 dataset. See annotation guidelines in Appendix A. In addition, annotators revise ASR data to generate subtitles.
Character photos. Annotators identify main characters and select a face for each in each movie clip, creating a cast list with names and corresponding faces (see Figure 3).
Audio-visual character identification. Annotators identify the speaker for each dialogue line in the clips. If the speaker is in the cast list, their character name is assigned; otherwise, a descriptive name is used. Figure 3 provides an example of audio-visual character identification.
4.3 Evaluation of Description
Evaluating long video descriptions is challenging due to the numerous details they contain, which increases evaluation costs. To address this, we develop an automatic evaluation method called StoryQA for the MovieStory101 test set. On average, each 3-minute video in StoryQA is accompanied by 38 multiple-choice questions focusing on visual actions, character relationships, and plot development. During evaluation, we employ GPT-4 [31] to answer these questions given the long description as context, and the QA accuracy serves as an automated metric for assessing the quality of the description. The prompt template is shown in Appendix F. Table 3 shows examples in StoryQA.
Figure 5: The distribution of the three types of QA counts across each clip in the MovieStory101 test set.
Creation for StoryQA We identify three key types of information for evaluating long video descriptions: 1) Action: actions, their sequence, and causal relationships; 2) Character: roles, relationships, and their evolution; 3) Plot: causes, character motivations, and event connections. We generate QA pairs for each type, as detailed in Appendix B. Figure 5 shows the distribution of QA counts across each clip in the MovieStory101 test set.
Validation for StoryQA To evaluate the quality of StoryQA, we sample 100 QA pairs, checking if each question is answerable based on the corresponding video and if the correct answer is among the options. The qualification rate is 98%. Then, we evaluate the GPT-4 evaluator by sampling 100 (description, question) pairs, with descriptions from both baselines in the Experiments section and StoryTeller. The authors answer each question, and we measure GPT-4’s agreement with the authors, which is 95%.
5 Experiments
5.1 Baselines
We compare our method against two types of baselines:
Model StoryQA Accuracy
@@ 5.1 Baselines
Table 4: The results of video descriptions generated by StoryTeller and baselines evluated on StoryQA.
@@
For all baselines, we provide the video, cast list and subtitles as input. Specifically, for Gemini-1.5-pro, we input the entire 3-minute video, while for other models, we divide the video into 10-second clips and sample 8 frames from each. Finally, we concatenate the descriptions from each clip to form a complete description of the entire 3-minute video.
@@
5.2 Experimental Results
First, we evaluate the descriptions generated by our method and the baselines on the StoryQA dataset. As shown in Table 4, StoryTeller outperforms all baselines on StoryQA. Notably, StoryTeller achieves a 9.5% higher accuracy than the strongest baseline, Gemini-1.5-pro. Table 5 shows a video example featuring multiple characters, where StoryTeller effectively tracks their relationships, resulting in description with fewer hallucination compared to Gemini-15-pro. When analyzing performance across different question types in StoryQA, StoryTeller outperforms all baselines on action, character and plot questions, indicating that StoryTeller provides more accurate descriptions of both low-level visual details and high-level plot information.
@@ 5.2 Experimental Results
Table 7: Evaluation of StoryTeller as a general framework for improving video descriptions across vairous LVLMs. The performance on StoryQA when key components are incrementally added: video segmentation module, integration of the audio-visual character identification (with ‘Others’ label for unlisted characters), and the use of descriptive names for unlisted characters. The baseline approach involves segmenting videos into 10-second clips and generating descriptions without audio-visual character identification.
@@
The system design of StoryTeller is a general framework applicable to various LVLMs to enhance their long video descriptions. Table 7 presents the additive contributions of each component for various LVLMs. First, replacing the direct division into 10-second clips with our video segmentation module consistently improves description accuracy across all models, resulting in an average relative increase of 6.3% in accuracy on the StoryQA. Second, utilizing the audio-visual identification results generated by our model as input for the LVLMs, we observe an average relative accuracy improvement of 6.4%. This demonstrates that our model’s audio-visual identification results is effective for enhancing video descriptions across various LVLMs. Lastly, we observe that replacing the label ‘Others’ with descriptive names generated by our model further improves video descriptions, consistently benefiting all models. Notably, under our framework, Gemini-1.5-pro and GPT-4o achieve relative improvements of 5.5% and 13.0% in accuracy on StoryQA, respectively.
@@
At last, we conduct ablation study to evaluate the necessity of multi-modal inputs, fine-tuning for audio and visual tasks (audio diarization, character recognition, identification, and tracking), as well as effect of global decoding within the audio-visual identification module. We assess the accuracy of audio-visual character identification on the MovieStory101 test set, as shown in Table 8. It indicates that combining multi-modal inputs is crucial for character identification. Specifically, removing subtitle, audio or video inputs results in a reduction in accuracy by 5.0%, 6.9% and 36.1%, respectively. In addition, fine-tuning the audio and visual tasks improves performance by 4.6%, while global decoding contributes 3.2% improvement.